Identify the electric flux passing through a rectangle with sides of 46 m and 98 m found in a region with a uniform electric field of 700 N/C and an angle of 80º.


1
Expert's answer
2022-02-07T09:55:08-0500

Given:

"A=\rm 46\: m\times 98\: m=4508\: m^2"

"E=700\:\rm N/C"

"\theta =80^\circ"


The electric flux passing through a rectangle

"\phi=EA\cos\theta\\
=700*4508*\cos 80^\circ=5.5*10^5\: \rm N\cdot m^2/C"